# Neural Concept Binder

### v1

@@ 3.3 Inspecting NCB’s Concepts
Finally, (iv) Similarity inspection allows inspecting NCB’s continuous encoding space on a more global level (in comparison to the more symbolic, sample-based inspection above), e.g., "What are similar concepts to this concept?". Specifically, NCB’s distance metric $d$ directly provides information about the similarity between concepts in the continuous representation. Inspecting the block-slot encoding space thus allows to identify a suboptimal soft binding, e.g., when block encodings are similar according to $g$ but not according to the human stakeholder. Overall, these inspection mechanisms allow a human stakeholder to ask a diverse set of questions concerning a model’s learned concepts.
3.4 Revising NCB’s Concepts
Let us now describe how a human stakeholder can revise NCB’s concept space. Below, we provide details on the three main actions for symbolic revision (i.e., revision on the representations in $\mathcal{R}$): (i) merging, (ii) deleting, or (iii) adding information. These actions can be performed on a single encoding or on a concept level and essentially represent a form of "reorganization" of information stored in $\mathcal{R}$. Furthermore, we provide details on how to (iv) revise the continuous latent space, which essentially requires finetuning of $g$’s parameters.
(i) Merge Concepts: In the case that $\mathcal{R}$ contains multiple concepts that, according to human knowledge, represent a joint underlying concept (e.g., two concepts for purple in Fig.  3 (right)) it is easy to update the model’s internal representations by replacing the concept symbols of one concept with those of the second concept. Formally, (in the case of two concepts) if $v_{l}$ and $v_{m}$ of block $j$ should be merged, the information stored in $\mathcal{R}^{j}$ is updated via $\forall(\texttt{enc}_{l}^{j},v_{l}):v_{l}\rightarrow v_{m}$.
(ii) Delete Encodings or Concepts: If $\mathcal{R}$ contains an encoding for a specific concept that does not match the other encodings of that concept, e.g., a misplaced exemplar, this encoding can simply be removed via $\texttt{remove}((\texttt{enc}_{l}^{j},v_{l}))$. Accordingly, if an entire concept is identified as suboptimal, it is possible to delete all corresponding encodings of concept $m$ in $\mathcal{R}^{j}$ as $\forall(\texttt{enc}_{l}^{j},v_{l}):v_{l}=m\implies\texttt{remove}((\texttt{enc}_{l}^{j},v_{l}))$.
(iii) Add Encodings or Concepts: In a similar fashion, it is also possible to add encodings to $\mathcal{R}$. For instance, if a specific concept is not sufficiently well captured via the encodings in $\mathcal{R}^{j}$, one can add an additional encoding, $\hat{\texttt{enc}}_{l}^{j}$, for the concept, $m$, via $\texttt{add}((\hat{\texttt{enc}}_{l}^{j},m))$. Accordingly, it is also possible to add encodings for an entire concept, $\hat{v}_{l}$: $\forall(\hat{\texttt{enc}}_{l}^{j},\hat{v}_{l})\implies\texttt{add}((\texttt{enc}_{l}^{j},m))$
(iv) Revise the (Continuous) Latent Space: Lastly, if the soft binder provides suboptimal object- and factor-level block-slot encodings, it is further possible to integrate revisory feedback on the soft binder’s continuous latent space. This can be achieved via additional finetuning of the soft binder’s parameters, $\theta$, e.g., via standard forms of weak supervision  [35, 60] or interactive learning [59, 52].
@@ 3.4 Revising NCB’s Concepts
Table 1: Comparison of different approaches for concept learning. Hereby, we differentiate based on the following categories: whether a method (1) is learned in an unsupervised fashion, (2) provides object-level concepts (i.e., can explicitly process multiple objects), (3) provides factor-level concepts (i.e., the color green), (4) provides continuous concept encodings, (5) provides discrete concept encodings, (6) provides inherently inspectable and (7) revisable concept representations.
@@
In summary, our novel Neural Concept Binder framework fulfills several important desiderata for concept learning (cf. Tab.  1). Specifically, NCB learns concepts in an unsupervised fashion that are structured on both an object (i.e., can explicitly process multiple objects) and factor-level. Furthermore, next to standard continuous encodings, NCB also provides discrete concept representations, which are crucial for interpretability and integration into symbolic computations. Lastly, NCB’s concept space is inspectable and revisable, essential for unsupervised learned concept representations.
@@ 4 Experimental Evaluations
CLEVR-Sudoku. To investigate the potential of integrating NCB’s discrete concept representations into symbolic downstream tasks, we introduce the novel CLEVR-Sudoku dataset. It represents a challenging visual puzzle requiring both visual object perception and reasoning capabilities. Each sample of the dataset (cf. Fig.  4 for an example puzzle) consists of a Sudoku puzzle that is (partially filled) with CLEVR-based images [24] and additional example images that depict the mapping of relevant object properties to digits. Specifically, each digit of the Sudoku is replaced by an image of an object. All objects representing the same digit share a set of common properties, e.g., in Fig.  4 all objects replacing "1"s are yellow spheres. We introduce two variants of CLEVR-Sudoku: Sudoku CLEVR-Easy and Sudoku CLEVR. In the first one, all objects are large and metallic, leaving only shape and color as properties for the digits. In Sudoku CLEVR, the size and material of the CLEVR objects are also relevant for the digits. Moreover, up to 10 example images are provided per digit mapping; the fewer examples provided, the harder it
@@
is to learn the mapping. The initial state and digit-attribute mapping vary for each sample. One specific intricacy of CLEVR-Sudoku is that a Sudoku can only be solved if all subcell images are correctly mapped to the digits. Even one mistake can lead to an unsolvable Sudoku. Thus, in comparison to standard Sudoku puzzles, which mainly require deductive reasoning, solving CLEVR-Sudoku additionally requires complex object recognition and the ability to map these visual concept perceptions to the task concepts (i.e. the 9 digits of Sudoku). We refer to Suppl. B for further details.
@@
Models. For our evaluations, we instantiate Neural Concept Binder based on the SysBinder model [56] for the soft binder encoder, $g$, and hdbscan [8, 9] for the clustering model, $h$. We refer to Sec. A.3 for all additional details. In the context of (Q1), we compare NCB’s results to those of four different variations of the SysBinder model [56]. We denote the original SysBinder configuration as SysBinder (cont.) (provides continuous block-slot encodings). For SysBinder, we discretize SysBinder’s continuous encodings at inference time (via an $\operatorname*{argmin}$ operation over SysBinder’s internal codebooks). SysBinder (hard) is trained from the start to provide discrete encodings (via a low codebook softmax temperature). SysBinder (step) is trained by step-wise decreasing this temperature (cf. Suppl. D for details). For evaluations on CLEVR-Sudoku (Q2+Q3), we first infer NCB’s discrete concept-slot encodings for a puzzle’s candidate examples. These encodings (together with the corresponding digit labels) are passed to a symbolic classifier, which is fit to predict the digits from the encodings. Next, this classifier infers the digits from the image of each subcell in the puzzle’s initial state. These are used by a constraint propagation and search-based algorithm [46, 6] to solve the puzzle (cf. Sec. E.2 for details). We denote the combination of the symbolic classifier and constraint solver as solver. We compare the solver’s performance when provided ground-truth (GT) object-property labels (GT concepts), encodings from a supervised trained slot attention encoder [36] (SA (supervised)) and the discrete encodings from SysBinder (denoted here as SysBinder (unsupervised)). For classification evaluations (Q4), we evaluate a configuration in which a set transformer classifier [31] is provided NCB’s concept encodings (NCB + NN) and makes a final class prediction based on these (cf. Sec. E.4). We compare to SA + NN, i.e., where a supervised trained slot attention encoder [36] provides object-property predictions.
@@
Metrics. We evaluate all models based on their accuracies on held-out test splits, each with 3 seeded reruns. We provide average accuracies and standard deviations over these. When assessing the expressiveness of NCB’s concept-slot encodings (Q1), we evaluate the accuracy for object-property prediction. When evaluating the performance of the downstream tasks, we provide the percentage of solved CLEVR-Sudokus (Q2) and the classification accuracy on the test set of CLEVR-Hans3 (Q4).
@@
Dataset N Train SysBinder ((((cont.)))) SysBinder SysBinder (hard) SysBinder (step) Neural Concept Binder
CLEVR- Easy N=2000 $\mathbf{\bullet\ 99.83}\mbox{\scriptsize$\pm 0.24$}$ $92.49\mbox{\scriptsize$\pm 5.45$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $95.76\mbox{\scriptsize$\pm 4.92$}$ $\mathbf{\circ\ 99.02}\mbox{\scriptsize$\pm 1.00$}$
N=200 $\mathbf{\bullet\ 99.20}\mbox{\scriptsize$\pm 0.41$}$ $87.90\mbox{\scriptsize$\pm 8.05$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $92.42\mbox{\scriptsize$\pm 7.32$}$ $\mathbf{\circ\ 98.50}\mbox{\scriptsize$\pm 1.80$}$
N=50 $\mathbf{\circ\ 91.13}\mbox{\scriptsize$\pm 4.21$}$ $78.41\mbox{\scriptsize$\pm 8.69$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $70.64\mbox{\scriptsize$\pm 11.89$}$ $\mathbf{\bullet\ 95.87}\mbox{\scriptsize$\pm 2.93$}$
N=20 $\mathbf{\circ\ 64.88}\mbox{\scriptsize$\pm 10.89$}$ $62.61\mbox{\scriptsize$\pm 7.18$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $54.61\mbox{\scriptsize$\pm 9.57$}$ $\mathbf{\bullet\ 94.22}\mbox{\scriptsize$\pm 4.11$}$
CLEVR N=2000 $\mathbf{\bullet\ 98.86}\mbox{\scriptsize$\pm 1.15$}$ $86.22\mbox{\scriptsize$\pm 10.40$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $88.90\mbox{\scriptsize$\pm 14.81$}$ $\mathbf{\circ\ 97.26}\mbox{\scriptsize$\pm 2.67$}$
N=200 $\mathbf{\bullet\ 97.61}\mbox{\scriptsize$\pm 2.58$}$ $81.13\mbox{\scriptsize$\pm 12.39$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $83.17\mbox{\scriptsize$\pm 17.05$}$ $\mathbf{\circ\ 96.80}\mbox{\scriptsize$\pm 3.01$}$
N=50 $\mathbf{\circ\ 93.25}\mbox{\scriptsize$\pm 4.62$}$ $61.67\mbox{\scriptsize$\pm 8.51$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $68.81\mbox{\scriptsize$\pm 17.74$}$ $\mathbf{\bullet\ 94.67}\mbox{\scriptsize$\pm 4.65$}$
N=20 $\mathbf{\circ\ 79.11}\mbox{\scriptsize$\pm 8.75$}$ $49.79\mbox{\scriptsize$\pm 6.73$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $58.58\mbox{\scriptsize$\pm 16.09$}$ $\mathbf{\bullet\ 88.57}\mbox{\scriptsize$\pm 4.68$}$
@@ 4.1 Evaluations
Discrete, yet expressive representations (Q1). In (Q1), we investigate how much valuable information NCB’s discrete concept-slot encodings contain, despite NCB’s inherent information bottleneck. For this purpose, we train a classifier on NCB’s encodings to predict corresponding object-property labels, e.g., the color green (cf. Sec. E.1 for details). In Tab.  2 we provide results for the CLEVR-Easy and CLEVR dataset with a classification training set of 2000, 200, 50, or 20 encodings. Focusing first on the results for $N=2000$, we observe that, as expected, the continuous representation of the original Sysbinder model contains the most information compared to all discrete encodings. Remarkably, however, NCB’s discrete concept representations are nearly on par with the continuous encodings. This is particularly remarkable given the immense information bottleneck in NCB  [^2] SysBinder (cont.) provides $2048$-sized continuous encodings, whereas NCB provides discrete encodings of size $\leq 16$.. Moreover, we observe that NCB’s encodings greatly outperform all other forms of discrete representations. Focusing now on the results when the classifier is trained on data subsets, we observe a substantial degradation in performance via encodings of any of the discrete baselines, and the continuous encodings. In stark contrast, when classifying based on NCB’s encodings, the accuracy remains nearly constant even with 1% of the initial training samples. We provide additional ablations on the effect of concept encoding types and NCB’s selection function in Sec. F.1 as well as an ablation analysis on the effect of suboptimal behavior from NCB’s individual components in Sec. F.2. We further refer to Sec. F.3 for an analysis of NCB’s obtained concept space as well as Suppl. G for examples of learned concepts. Overall, our results provide evidence for the expressiveness of NCB’s concept encodings despite their enormous information bottleneck. Furthermore, our results for $N\leq 200$ suggest that the information stored in the NCB’s encodings is easier to generalize compared to the baselines. We thus answer Q1 affirmatively.
@@
Utilising unsupervised concepts for solving visual Sudoku (Q2). In our following evaluations, we investigate the potential of NCB’s representations for solving complex reasoning tasks via their integration into symbolic computations. We base these evaluations on our novel CLEVR-Sudoku dataset. We report the percentage of solved puzzles for CLEVR-Sudoku in Fig.  5. Hereby, it is important to note that the solver can only solve a puzzle if each of the images of the initial state has been classified correctly. Thus, the results of Fig.  5 represent "all-or-nothing" results. Focusing on the results left of the dashed lines, we observe that the symbolic module can solve every puzzle with GT concepts, even when only one example image is provided. Interestingly, performance drops drastically when provided with encodings of SA (supervised). This highlights the difficulty of the CLEVR-Sudoku puzzles: minor errors in digit prediction can lead to major errors in solving the underlying puzzle. When comparing the performance via encodings from the two unsupervised models, we observe quite strong performances via NCB’s concept encodings. e.g., they allow to solve $\approx 50\%$ for the 10 example configurations (over both data settings) in comparison to $\approx 61\%$ for SA (supervised). In contrast, when provided SysBinder’s encodings the solver fails over all Sudoku variations. This demonstrates the effectiveness of NCB’s binding mechanisms over solely those of the SysBinder approach. We refer to Sec. F.4 for further discussions as well as quantitative digit classification results. Overall, our evaluations highlight the potential of NCB’s unsupervised concept encodings for solving complex, symbolic downstream tasks. We therefore answer Q2 affirmatively.
@@
Easily revising NCB’s concepts (Q3). In our next evaluations, we illustrate the potential of NCB’s revision procedures. As revision on the continuous latent space of NCB’s soft binder is analogous to existing approaches (e.g., [52, 49, 60]), we focus these evaluations on the novel, NCB-specific forms of symbolic revision, i.e., on the hard binder’s concept space. We hereby illustrate the first two forms of symbolic revision (removing and merging concept information) via two forms of feedback: based on a pretrained vision-language model (here via GPT-4 [47]) and simulated human feedback. In both cases, we ask the revisory agent to identify which concepts in each block should be removed or merged based on exemplar images of each concept, i.e., implicit concept inspection (cf. Sec. E.3 for details). In Fig.  5 we provide CLEVR-Sudoku performances when the NCB’s retrieval corpus has been updated based on the different revisory agents (i.e., NCB revised (GPT-4) and NCB revised (human)). Interestingly, although we observe improvements in the few example settings via GPT-4’s revisions, we observe a negative effect in the settings with more digit examples. This is due to the suboptimal behavior of GPT-4 in providing consistent object descriptions, such that too much concept information is removed or merged and highlights the effect of “ill-informed” feedback (cf. Sec. F.5). In contrast, when provided with human revision, we observe a substantial boost in Sudoku performance, particularly in puzzle configurations with few candidate examples. Moreover, via NCB’s similarity inspection mechanism (cf. Sec.  3.3), a human stakeholder can easily identify models that provide suboptimal soft binding processing. In this case, such models can be excluded from further downstream evaluations (cf. NCB revised (human)*) and must be refined via finetuning $g$’s parameters (e.g., via approaches of [52, 49, 60]). In Sec. F.6, we further investigate concept revision via adding information. Overall, our results highlight the potential and ease of revising NCB’s concept space. We thus answer Q3 positively.

### v2

@@ 3.3 Inspecting and Revising NCB’s Concepts
Inspection. NCB inherently enables: (i) implicit, (ii) comparative, (iii) interventional and (iv) similarity-based inspection (cf. Fig.  3). Where the first three aim at investigating NCB’s explicit, symbolic concept space (stored in $\mathcal{R}$), the last one aims at investigating its latent, continuous concept space (stored in $\theta$). (i) Implicit inspection queries the model to provide a set of examples for a specific concept. Essentially, this answers the question "What are examples of this concept?". NCB answers this question in two ways: by providing samples from the retrieval corpus corresponding to exemplars of the concept or by identifying additional data samples belonging to the concept at hand. (ii) Comparative inspection, on the other hand, allows comparing two specifically different concepts, e.g., "Why does this object depict concept H5 and not concept H1?". NCB hereby provides examples for both concepts for the user to compare and potentially identify dissimilar properties. Ultimately, this form of inspection allows to answer questions of the form "Why not …?" and represents a valuable tool for in-depth and targeted concept inspection. (iii) Interventional inspection allows to answer questions such as "What if this object would have concept H1?" To answer this question, NCB utilizes its decoder $g^{\prime}$. Specifically, by swapping the block $z_{i}^{j}$ of a data sample’s block-slot encoding with that of a representative sample, $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$, NCB can provide an interventional image reconstruction, from which the effect of the swapped concept can be observed. Ultimately, this form of inspection allows to answer important questions of the form "What if …?". Finally, (iv) Similarity inspection allows inspecting NCB’s continuous encoding space on a more global level (in comparison to the more symbolic, sample-based inspection above), e.g., "What are similar concepts to this concept?". Specifically, NCB’s distance metric $d$ directly provides information about the similarity between concepts in the continuous representation. Inspecting the block-slot encoding space thus allows to identify a suboptimal soft binding, e.g., when block encodings are similar according to $g$ but not according to the human stakeholder. Overall, these inspection mechanisms allow a human stakeholder to ask a diverse set of questions concerning a model’s learned concepts (cf. Fig. 15, Fig. 16 and Fig. 17 for additional examples of the inspection types).
Revise. Let us now describe how a human stakeholder can revise NCB’s concept space. Below, we provide details on the three main actions for symbolic revision (i.e., revision on the representations in $\mathcal{R}$): (i) merging, (ii) deleting, or (iii) adding information. These actions can be performed on a single encoding or on a concept level and essentially represent a form of "reorganization" of information stored in $\mathcal{R}$. Furthermore, we provide details on how to (iv) revise the continuous latent space, which essentially requires finetuning of $g$’s parameters. (i) Merge Concepts: In the case that $\mathcal{R}$ contains multiple concepts that, according to additional knowledge (e.g., from a human or other model), represent a joint underlying concept (e.g., two concepts for purple in Fig.  3 (right)) it is easy to update the model’s internal representations by replacing the concept symbols of one concept with those of the second concept. Specifically, for block $j$ if concept $m$ should be merged with concept $b$ where $m,b\in\{1,\cdots,N_{C}\}$, then for all corpus tuples, $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$, we replace $v_{l}$ with $b$ if $v_{l}=m$. (ii) Delete Encodings or Concepts: If $\mathcal{R}^{j}$ contains an encoding, $\texttt{enc}_{l}^{j}$, for a specific concept, $m$, that does not match the other encodings of that concept (e.g., a misplaced exemplar) this encoding can simply be deleted from the corpus. Accordingly, if an entire concept, $m$, is identified as suboptimal, one can simply delete all corresponding encodings of that concept. I.e., for all corpus tuples, $(\texttt{enc}_{l}^{j},v_{l})\in\mathcal{R}^{j}$, we remove the tuple if $v_{l}=m$. (iii) Add Encodings or Concepts: If a specific concept is not sufficiently well captured via the existing encodings in $\mathcal{R}^{j}$, one can simply add a new encoding, $\hat{\texttt{enc}}_{l+1}^{j}$, for the concept, $m$, to the corpus. This leads to an additional entry in the corpus, $(\hat{\texttt{enc}}_{l+1}^{j},m)$. Accordingly, it is also possible to add encodings for an entire concept. Hereby, one gathers block encodings of objects that represent that novel concept and adds these to the corpus as $(\hat{\texttt{enc}}_{l+1}^{j},b)$ with $b=N_{C}+1$. (iv) Revise the (Continuous) Latent Space: Lastly, if the soft binder provides suboptimal object- and factor-level block-slot encodings, it is further possible to integrate revisory feedback on the soft binder’s continuous latent space. This can be achieved via additional finetuning of the soft binder’s parameters, $\theta$, e.g., via standard forms of weak supervision  [42, 69] or interactive learning [68, 61].
@@
Table 1: Comparison of different approaches for concept learning. Hereby, we differentiate based on the following categories: whether a method (1) is learned in an unsupervised fashion, (2) provides object-level concepts (i.e., can explicitly process multiple objects), (3) provides factor-level concepts (e.g., the color green), (4) provides continuous concept encodings, (5) provides discrete concept encodings, (6) provides inherently inspectable and (7) revisable concept representations.
@@
In summary, our novel Neural Concept Binder framework fulfills several important desiderata for concept learning (cf. Tab.  1). Specifically, NCB learns concepts in an unsupervised fashion that are structured on both an object and factor-level. Furthermore, next to standard continuous encodings, NCB also provides discrete concept representations, which are crucial for interpretability and integration into symbolic computations. Lastly, NCB’s concept space is inspectable and revisable, essential for unsupervised learned concept representations.
@@ 4 Experimental Evaluations
CLEVR-Sudoku. To investigate the potential of integrating NCB’s discrete concept representations into symbolic downstream tasks, we introduce the novel CLEVR-Sudoku dataset. This dataset presents a challenging visual puzzle that requires both visual object perception and reasoning capabilities. Each sample in the dataset (cf. Fig.  4 for an example puzzle) consists of a Sudoku puzzle (partially filled) with CLEVR-based images [30] and additional example images depicting the mapping of relevant object properties to digits. Specifically, each digit in the Sudoku is replaced by an image of an object. All objects representing the same digit share a set of common properties, e.g., in Fig.  4, all objects replacing "1"s are yellow spheres.
@@
We introduce two variants of CLEVR-Sudoku: Sudoku CLEVR-Easy and Sudoku CLEVR. In the first variant, shape and color are distinguishing properties for the digits. In Sudoku CLEVR, additional object attributes — size and material — are relevant for the digit identification. Moreover, up to 10 example images are provided per digit mapping; the fewer examples provided, the more difficult it becomes to learn the mapping. The initial state and digit-attribute mappings vary across samples. One specific intricacy of CLEVR-Sudoku is that the puzzle can only be solved if all subcell images are correctly mapped to their corresponding digits. Even a single mistake can render the Sudoku unsolvable. Thus, compared to standard Sudoku puzzles, which primarily require deductive reasoning, solving CLEVR-Sudoku also demands complex object recognition and the ability to map visual concept perceptions to the task concepts (i.e., the 9 digits of Sudoku). For further details, we refer to Suppl. B.
@@
Models. For our evaluations, we instantiate Neural Concept Binder based on the SysBinder model [65] for the soft binder encoder, $g$, and HDBSCAN [12, 13] for the clustering model, $h$. Further details about the instantiation can be found in Sec. A.4. In the context of Q1, we compare NCB’s results to four variations of the SysBinder model [65], as well as the recent Neural Language of Thought Model (NLOTM) [80]. We refer to the original SysBinder configuration as SysBinder (cont.), which provides continuous block-slot encodings. In SysBinder, SysBinder’s continuous encodings are discretized at inference time via an $\operatorname*{argmin}$ operation over its internal codebooks. SysBinder (hard) is trained from the beginning to produce discrete encodings using a low codebook softmax temperature. SysBinder (step) is trained with a step-wise decrease in temperature (cf. Suppl. D for details). For evaluations on CLEVR-Sudoku (Q2 and Q3), we first infer NCB’s discrete concept-slot encodings from the puzzle’s candidate examples. These encodings, along with their corresponding digit labels, are then passed to a symbolic classifier, which is trained to predict digits from the encodings. The classifier subsequently infers the digits for each subcell in the puzzle’s initial state. These predictions are used by a constraint propagation and search-based algorithm [55, 10] to solve the puzzle (cf. Sec. E.2 for details). We refer to the combination of the symbolic classifier and constraint solver as the solver. We compare the solver’s performance when provided with ground-truth (GT) object-property labels (GT concepts), encodings from a supervised slot attention encoder [43] (SA (supervised)), and the discrete encodings from SysBinder (denoted as SysBinder (unsupervised)). For classification evaluations (Q4), we evaluate a configuration in which a set transformer classifier [37] is provided with NCB’s concept encodings (NCB + NN) to make final class predictions (cf. Sec. E.4). We compare this to SA + NN, where a supervised slot attention encoder [43] provides object-property predictions.
@@
Metrics. We evaluate all models based on their accuracies on held-out test splits, each with 3 seeded runs. We provide average accuracies and standard deviations over these. When assessing the expressiveness of NCB’s concept-slot encodings (Q1), we evaluate the accuracy for object-property prediction. When evaluating the performance of the downstream tasks, we provide the percentage of solved CLEVR-Sudokus (Q2) and the classification accuracy on the test set of CLEVR-Hans3 (Q4).
@@
Dataset N Train SysBinder ((((cont.)))) SysBinder SysBinder (hard) SysBinder (step) NLOTM Neural Concept Binder
CLEVR- Easy N=2000 $\mathbf{\bullet\ 99.83}\mbox{\scriptsize$\pm 0.24$}$ $92.49\mbox{\scriptsize$\pm 5.45$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $95.76\mbox{\scriptsize$\pm 4.92$}$ $84.36\mbox{\scriptsize$\pm 8.54$}$ $\mathbf{\circ\ 99.02}\mbox{\scriptsize$\pm 1.00$}$
N=200 $\mathbf{\bullet\ 99.20}\mbox{\scriptsize$\pm 0.41$}$ $87.90\mbox{\scriptsize$\pm 8.05$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $92.42\mbox{\scriptsize$\pm 7.32$}$ $72.99\mbox{\scriptsize$\pm 8.43$}$ $\mathbf{\circ\ 98.50}\mbox{\scriptsize$\pm 1.80$}$
N=50 $\mathbf{\circ\ 91.13}\mbox{\scriptsize$\pm 4.21$}$ $78.41\mbox{\scriptsize$\pm 8.69$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $70.64\mbox{\scriptsize$\pm 11.89$}$ $49.94\mbox{\scriptsize$\pm 4.97$}$ $\mathbf{\bullet\ 95.87}\mbox{\scriptsize$\pm 2.93$}$
N=20 $\mathbf{\circ\ 64.88}\mbox{\scriptsize$\pm 10.89$}$ $62.61\mbox{\scriptsize$\pm 7.18$}$ $22.92\mbox{\scriptsize$\pm 0.00$}$ $54.61\mbox{\scriptsize$\pm 9.57$}$ $37.05\mbox{\scriptsize$\pm 4.11$}$ $\mathbf{\bullet\ 94.22}\mbox{\scriptsize$\pm 4.11$}$
CLEVR N=2000 $\mathbf{\bullet\ 98.86}\mbox{\scriptsize$\pm 1.15$}$ $86.22\mbox{\scriptsize$\pm 10.40$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $88.90\mbox{\scriptsize$\pm 14.81$}$ $54.10\mbox{\scriptsize$\pm 18.78$}$ $\mathbf{\circ\ 97.26}\mbox{\scriptsize$\pm 2.67$}$
N=200 $\mathbf{\bullet\ 97.61}\mbox{\scriptsize$\pm 2.58$}$ $81.13\mbox{\scriptsize$\pm 12.39$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $83.17\mbox{\scriptsize$\pm 17.05$}$ $50.17\mbox{\scriptsize$\pm 16.26$}$ $\mathbf{\circ\ 96.80}\mbox{\scriptsize$\pm 3.01$}$
N=50 $\mathbf{\circ\ 93.25}\mbox{\scriptsize$\pm 4.62$}$ $61.67\mbox{\scriptsize$\pm 8.51$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $68.81\mbox{\scriptsize$\pm 17.74$}$ $43.60\mbox{\scriptsize$\pm 13.38$}$ $\mathbf{\bullet\ 94.67}\mbox{\scriptsize$\pm 4.65$}$
N=20 $\mathbf{\circ\ 79.11}\mbox{\scriptsize$\pm 8.75$}$ $49.79\mbox{\scriptsize$\pm 6.73$}$ $36.46\mbox{\scriptsize$\pm 0.00$}$ $58.58\mbox{\scriptsize$\pm 16.09$}$ $41.52\mbox{\scriptsize$\pm 12.90$}$ $\mathbf{\bullet\ 88.57}\mbox{\scriptsize$\pm 4.68$}$
@@ 4.1 Evaluations
Discrete, yet expressive representations (Q1). First, we investigate how much valuable information NCB’s discrete concept-slot encodings contain, despite NCB’s inherent information bottleneck. To assess this, we train a classifier on NCB’s encodings to predict corresponding object-property labels, e.g., the color green (cf. Sec. E.1 for details). In Tab.  2, we present the results for the CLEVR-Easy and CLEVR datasets, using classification training sets with 2000, 200, 50, or 20 encodings. Focusing first on the results for $N=2000$, we observe that, as expected, the continuous representation of the original SysBinder model contains more information compared to all discrete encodings. Remarkably, however, NCB’s discrete concept representations are nearly on par with the continuous encodings. This is particularly notable given NCB’s immense information bottleneck [^1] SysBinder (cont.) provides 2048-sized continuous encodings, whereas NCB provides discrete encodings of size $\leq 16$.. Additionally, we observe that NCB’s encodings significantly outperform all other forms of discrete representations. Shifting focus to the results when the classifier is trained on data subsets, we observe a substantial degradation in performance when using encodings from any of the discrete baselines or the continuous encodings. In stark contrast, when classifying based on NCB’s encodings, the accuracy remains nearly constant, even with just 1% of the initial training samples. We provide additional ablations on the effect of concept encoding types and NCB’s selection function in Sec. F.1, as well as an ablation analysis on the effect of suboptimal behavior from NCB’s individual components in Sec. F.2. Further analysis of NCB’s concept space can be found in Sec. F.3, along with qualitative examples of learned concepts in Suppl. G. Overall, our results demonstrate the expressiveness of NCB’s concept encodings despite their significant information bottleneck. Furthermore, our results suggest that NCB’s encodings are easier to generalize compared to the baselines. Thus, we answer Q1 affirmatively.
@@
Utilizing unsupervised concepts for solving visual Sudoku (Q2). In our following evaluations, we investigate the potential of NCB’s representations for solving complex reasoning tasks through their integration into symbolic computations. These evaluations are based on our novel CLEVR-Sudoku dataset. The percentage of solved puzzles for CLEVR-Sudoku is reported in Fig.  5. It is important to note that the solver can only solve a puzzle if each image in the initial state has been classified correctly, meaning the results in Fig.  5 represent "all-or-nothing" outcomes. Focusing on the results to the left of the dashed lines, we observe that the symbolic module solves every puzzle with ground-truth (GT) concepts, even when only one example image is provided. Interestingly, performance drops significantly when using encodings from SA (supervised). This highlights the difficulty of the CLEVR-Sudoku puzzles: minor errors in digit prediction can lead to major failures in solving the puzzle. When comparing the performance of encodings from the two unsupervised models, we observe that NCB’s concept encodings perform quite well. E.g., they enable solving approximately 50% of the puzzles for the 10-example configurations, compared to approximately 61% for SA (supervised). In contrast, when using SysBinder’s encodings, the solver fails across all Sudoku variations. This demonstrates the effectiveness of NCB’s binding mechanisms over those of the SysBinder approach alone. We refer to Sec. F.4 for further discussions and quantitative digit classification results (Fig. 10). Overall, our evaluations highlight the potential of NCB’s unsupervised concept encodings for solving complex symbolic downstream tasks. We therefore answer Q2 affirmatively.
@@
Easily revising NCB’s concepts (Q3). In our next evaluations, we illustrate the potential of NCB’s revision procedures. Since revising the continuous latent space of NCB’s soft binder is analogous to existing approaches (e.g., [61, 58, 69]), we focus on the novel, NCB-specific forms of symbolic revision, i.e., revisions within the hard binder’s concept space. We demonstrate two forms of symbolic revision (removing and merging concept information) using feedback from two sources: a pretrained vision-language model (here via GPT-4 [56]) and simulated human feedback. In both cases, we ask the revisory agent to identify which concepts in each block should be removed or merged based on exemplar images of each concept, i.e., implicit concept inspection (cf. Sec. E.3 for details). In Fig.  5, we show CLEVR-Sudoku performance when NCB’s retrieval corpus is updated by different revisory agents (i.e., NCB revised (GPT-4) and NCB revised (human)). Interestingly, while GPT-4’s revisions improve performance in settings with few examples, they have a negative impact when more digit examples are present. This is due to GPT-4’s suboptimal consistency in object descriptions, leading to the removal or merging of too much concept information. This highlights the potential issue of "ill-informed" feedback (cf. Sec. F.5). In contrast, human revisions provide a substantial boost in Sudoku performance, particularly in puzzle configurations with fewer candidate examples. Moreover, using NCB’s similarity inspection mechanism (cf. Sec.  3.3), a human stakeholder can easily identify models that suffer from suboptimal soft binding processing. In such cases, these models can be excluded from further downstream evaluations (cf. NCB revised (human)*) and refined by finetuning $g$’s parameters (e.g., via approaches from [61, 58, 69]). In Sec. F.6, we further explore concept revision by adding new information. Overall, our results demonstrate the potential and ease of revising NCB’s concept space, allowing us to answer Q3 positively.